OpenStudy (anonymous):

a cell has an emf of 2.0 volt and an internal resistance of 0.1 ohm. what is the potentioal difference across its terminal when it is supplying a current of 5 A? what is the maximum current the cell can supply?

OpenStudy (souvik):

use this...V=E-Ir E=emf V=potential difference across its terminal I=current r=internal resistance

OpenStudy (anonymous):

thanks

OpenStudy (chenping):

Emf=V+Ir I=5A r=0.1ohm Emf=2V Thus, V=2-.5=1.5V

OpenStudy (anonymous):

The max current that can be supplied is given by I = E/r It is when there is no ext resistance in the circuit. It is called the short circuit current.